OpenStudy (anonymous):

Find the particular solution of the differential equation dy/dx = (x-2)e^(-2y) satisfying the initial condition y(2)=ln(2)

OpenStudy (johnweldon1993):

Hmm..well it's separable...so divide both sides by \(\large e^{-2y}\) \[\large e^{2y}\frac{dy}{dx} = (x - 2)\] Then integrate both sides

OpenStudy (anonymous):

but how does that satisfy the initial condition?

OpenStudy (anonymous):

I've gotten out -(ln|-x^2+4x|)/2 but where to go from here?

OpenStudy (johnweldon1993):

The initial condition is what we put in at the end to solve for the constant... \[\large \frac{1}{2} \log(x^2 - 4x + C_1)\] is what I get..remember the +C when integrating

OpenStudy (johnweldon1993):

And from there is where you apply the initial condition y(2) = ln(2) so \[\large \ln(2) = \frac{1}{2}\ln((2)^2 - 4(2) + C_1)\] \[\large \ln(2) = \frac{1}{2}\ln(-4 + C_1)\] \[\large 2\ln(2) = \ln(C_1 - 4)\] \[\large 4 = C_1 - 4\] \[\large C_1 = 8\] Now just make that your constant... \[\large y(x) = \frac{1}{2}\log(x^2 - 4x + 8)\] would be your final answer